# Adver-City: Open-Source Multi-Modal Dataset for Collaborative Perception Under Adverse Weather Conditions

### v1

@@ IV Benchmark
Adver-City Clear Weather $53.22$ $44.49$ $54.19$ $44.95$ $46.70$ $41.53$ $42.76$ $39.50$
Soft Rain $52.50$ $43.73$ $53.80$ $42.32$ $42.30$ $40.45$ $39.42$ $34.78$
Heavy Rain $54.38$ $42.05$ $49.42$ $44.71$ $43.06$ $38.72$ $34.11$ $38.03$
Fog $52.98$ $43.24$ $52.81$ $40.04$ $38.68$ $36.68$ $37.55$ $34.11$
Foggy Heavy Rain $54.31$ $42.97$ $54.42$ $38.76$ $37.83$ $35.77$ $40.40$ $34.18$
Glare $53.71$ $42.29$ $42.43$ $38.55$
OPV2V [16] Culver City $79.52$ $70.27$
TABLE IV: Benchmark results from 3D object detection models trained on OPV2V and tested with unseen data. All experiments used CoBEVT [22] models provided by the authors of HEAL [23]. PointPillars [38] was used as the backbone for Lidar, while Lift-Splat-Shoot [39] was used for camera with EfficientNet [40] as the image encoder. While Lidar experiments were more influenced by object density, Lidar + Camera experiments were more influenced by adverse weather conditions.
The average performance drop between Adver-City and Culver City ($31\%$ for both modalities) was due to differences in sensor configurations, positioning, and scenario locations. Despite this, clear patterns emerged: LiDAR-based detection was less impacted by adverse weather ($3.9\%$ variation between weather conditions, on average) than LiDAR + Camera ($7.2\%$ variation). On the other hand, LiDAR-based detection was more affected by higher object density ($10.6\%$ variation) than LiDAR + Camera ($3.0\%$ variation).
Despite CARLA’s limitations, performance noticeably dropped as weather conditions became more challenging, with a $19\%$ decline for the LiDAR + Camera model on sparse day scenarios. Moreover, the results highlight the diversity of Adver-City’s scenarios, with a performance gap of nearly $29\%$ for LiDAR-based detection and $27\%$ for LiDAR + Camera between the highest- and lowest-scoring scenarios.
Since camera-based 3D object detection was not viable with HEAL’s models, we performed 2D object detection with YOLOv8 [41] trained on the COCO dataset [42] to assess the impact of adverse weather on camera-based detection. The detection was done using only the front-facing camera of the ego vehicle. As shown in Table V, the model performed slightly below the baseline YOLO models on similar weather conditions, with results being impacted by nighttime by almost $26\%$.
Scenario Camera (mAP50) Baseline
Day Night (Day)
S D S D
Soft Rain $46.7$ $44.4$ $33.9$ $35.1$ $57.1$ [43]
Heavy Rain $42.1$ $41.8$ $21.5$ $27.7$ $46.0$ [43]
Fog $48.6$ $41.4$ $42.1$ $34.2$ $53.6$ [44]
Foggy Heavy Rain $38.7$ $40.9$ $34.4$ $26.7$
Glare $38.7$ $38.9$
TABLE V: Benchmark results from 2D object detection performed with YOLOv8 [41] on Adver-City’s adverse weather scenarios. S: sparse, D: dense. While glare was the most challenging condition for daytime scenarios, heavy rain was the toughest for nighttime.
Standing out from the 3D object detection benchmarks, object density played a lesser role in 2D object detection, with the median result being very similar both for sparse ($38.7$) and dense ($38.9$) scenarios, which is caused by different annotation strategies between the two tasks. While occlusion may heavily impact 3D detection, it is less prevalent in 2D detection because a completely occluded object will not have a corresponding 2D bounding box. Therefore, even when using 2D object detection techniques, it might be important to correlate different cameras – from the ego vehicle, the connected vehicles, and the RSUs – for improved detection. This is a new avenue of research that can be explored with Adver-City.
V Conclusions
@@
Our benchmarks showed that models trained without adverse weather scenes suffer a significant drop in performance when used for inference in these conditions. In the future, we hope to investigate how existing CP models trained on Adver-City perform across its diverse weather conditions, but also to improve upon their designs to address shortcomings that may arise from their usage under adverse weather.
@@
As an open-source dataset, Adver-City’s open code can be easily extended to include additional weather conditions such as snow or dust, new road configurations, and maps that leverage weather and density variations. Sensor configurations can also be adjusted to generate new data or enhance the realism of CARLA’s sensors. We hope its open-source nature encourages users to share their work, extending the benefits of collaboration beyond the confines of perception algorithms.

### v2

@@ 4 Benchmark
Adver-City Clear Weather $46.70$ $41.53$ $42.76$ $39.50$ $56.55$ $47.88$ $57.94$ $51.83$
Soft Rain $42.30$ $40.45$ $39.42$ $34.78$ $60.59$ $52.70$ $59.60$ $49.07$
Heavy Rain $43.06$ $38.72$ $34.11$ $38.03$ $63.18$ $49.01$ $56.49$ $51.83$
Fog $38.68$ $36.68$ $37.55$ $34.11$ $59.82$ $47.54$ $59.06$ $49.10$
Foggy Heavy Rain $37.83$ $35.77$ $40.40$ $34.18$ $59.25$ $48.80$ $60.44$ $50.59$
Glare $42.43$ $38.55$ $58.97$ $48.32$
OPV2V Culver City $70.27$
Table 4: Results for 3D object detection from the collaborative CoBEVT [42] model. The model trained on OPV2V [41] was used for inference on the entire Adver-City dataset to assess its robustness to challenging weather conditions, while the model trained on Adver-City was used only on its test set.
Regarding the results of the model trained on OPV2V, differences in scenario location, sensor configuration, and positioning caused a performance drop between the Culver City and Adver-City scenarios. While OPV2V was captured exclusively in urban scenarios with $64$-channel LiDARs and frontal cameras mounted on the vehicle’s hood, our dataset was captured in both rural and urban scenarios with $32$-channel LiDARs and frontal cameras mounted on the windshield.
As expected from a model trained without adverse weather, the best result occurred for the clear weather scenario during the day, with the lowest result for heavy rain during the night. Despite CARLA’s limitations, performance dropped noticeably as weather conditions became more challenging, with a gap of $12.59$ between the highest and lowest scoring scenarios. Increasing object density from sparse to dense decreased performance for most scenarios, indicating the effects of occlusion on the model’s detection.
When trained on Adver-City, however, the model’s results showed a smaller gap to Culver City, indicating that it adequately learned from our dataset’s distinct weather conditions and sensor configurations. Although the results for clear weather scenarios during the day improved, nighttime and adverse weather scenarios exhibited a significantly higher improvement, an effect of OPV2V’s scenarios exclusively portraying clear weather during the day and illustrating the importance of utilizing adverse weather data for training models.
The detection scores of the CoBEVT model trained on Adver-City over the entire test set were $58.30/52.44/38.90$ (AP@30/50/70). In Table 5, we compare it with the results of this model provided by [21] on other datasets, highlighting the impact that adverse weather conditions have on collaborative perception models.
LiDAR + Camera
AP@50
Adver-City 52.4
DAIR-V2X 63.8
OPV2V 88.5
Table 5: Results for 3D object detection for CoBEVT [42] in distinct collaborative perception datasets. Results for DAIR-V2X [45] and OPV2V [41] provided by [21].
5 Limitations
One of the main benefits of simulation comes from the amount of control over scenarios. It enables uncommon real-life conditions such as heavy fog and intense glare to be seamless and extensively experimented with. Moreover, simulation is more scalable and accessible than real-world data acquisition, which requires costly sensor equipment to be captured. However, such benefits do not come without downsides, the main one being a reduction in realism. Despite recent advances, current simulators are still not capable of accurately simulating complex real-world phenomena such as adverse weather conditions.
For that matter, we found it imperative to assess the level of realism of Adver-City, comparing it to real-world datasets with extensive adverse weather samples. Due to the scarcity of adverse weather conditions in real-world CP datasets, we selected three single-vehicle datasets for this comparison. The first is BDD100K [44], collected in the USA and featuring diverse weather conditions such as fog, rain, and snow, although only containing camera data. The second is NuScenes [5], one of the most used single-vehicle datasets to date, featuring multi-modal data collected in the USA and Singapore of clear weather and rainy scenes. Lastly, Boreas [4] is a multi-modal Canadian dataset containing rainy, snowy, and sunny scenarios.
We restricted our analysis to images of rainy scenarios during the day, since not all selected datasets have sufficient samples of other types of data such as LiDAR point clouds or foggy scenarios. Additionally, since all three single-vehicle datasets were collected in urban settings, we selected image samples from Adver-City’s urban scenarios (urban intersection and urban non-junction) in rainy conditions (soft rain and heavy rain) collected by the frontal camera from the ego vehicle or the CAVs. This criterion restricted the analysis to $4,443$ images from Adver-City, for which an equal number of images matching the criterion were randomly selected from the real-world datasets.
Inspired by [29], we evaluated the domain gap between synthetic and real data using t-SNE [37] and FID (Fréchet Inception Distance) [9]. We used Scikit-Learn’s [27] implementation of t-SNE calculating it from features computed by EfficientNet [35], and [31]’s implementation of FID.
The t-SNE plot, presented in Figure 5, illustrates the domain gap between each of the datasets analyzed, with similar samples grouped together. From the amount of overlap between samples in the plot, it can be noticed that samples from Adver-City have less in common with the real-world datasets than they have among themselves. Yet, many regions of the plot show overlap between Adver-City’s samples and samples from each of the real-world datasets. For a better visualization of the samples and common characteristics among distinct regions of the plot, see the Supplementary Material for a t-SNE plot with points replaced by their corresponding images.
Figure 5: t-SNE plot generated with Scikit-Learn [27] from features computed by EfficientNet [35] of each dataset’s rainy images. Adver-City’s clusters have lower overlap than the real-world datasets, yet still having minor overlap with them.
Even though the t-SNE plot shows that Adver-City has some level of similarity to the real-world datasets, it does not provide a quantitative evaluation of that similarity, for which the FID score brings a clearer metric. Table 6 shows the FID scores for the rainy images of each pair of datasets, showing that Adver-City’s FID score compared to both BDD100K ($147.1$) and NuScenes ($145.2$) is roughly $50\%$ greater than the score between the two of them ($95.6$). However, the BDD100K-Boreas’s score ($189.5$) is almost double the value of the BDD100K-NuScenes score. As can be observed in the Supplementary Material, Boreas’s distinctiveness can be attributed to the presence of multiple raindrops on the camera lens, a feature not observed in the other datasets. This distinctive feature makes Boreas more dissimilar to BDD100K and NuScenes than Adver-City, indicating that the synthetic dataset’s realism is within an acceptable margin.
Adver-City BDD100K NuScenes Boreas
Adver-City 00 $147.1$ $145.2$ $205.4$
BDD100K $147.1$ 00 $95.6$ $189.5$
NuScenes $145.2$ $95.6$ 00 $160.7$
Boreas $205.4$ $189.5$ $160.7$ 00
Table 6: FID scores [9] comparing rainy image samples between Adver-City and real-world single-vehicle datasets. A value of 00 indicates an equal distribution, while higher values indicate more distinct features. BDD100K [44] and NuScenes [5] have the most similar rainy samples, while Boreas [4] has the most distinct rainy scenes of all datasets analyzed.
Another way to assess the realism of Adver-City’s synthetic weather is to evaluate the FID score between its rainy and clear weather samples, then compare this score with that of real-world datasets. As presented in Table 7, this FID score for Adver-City is smaller than that of real-world datasets, which also suggests a lower realism. However, it is merely $17\%$ smaller than BDD100K’s score and $23\%$ smaller than NuScenes, indicating a fair level of realism.
Rain VS. Clear
Adver-City 60.7
BDD100K 72.9
NuScenes 79.4
Boreas 184.8
Table 7: FID scores between each dataset’s rainy and clear weather images. Adver-City’s lower score indicates the difference between its rainy and clear weather scenes is smaller than the difference found in real-world datasets. However, its gap is merely $17\%$ smaller than BDD100K’s gap.
Thus, our analysis shows that Adver-City is less realistic than real-world datasets, an expected feature of a synthetic dataset. However, we have shown that it is not significantly dissimilar from them, making it a valuable and viable source of experimentation for CP models focused on adverse weather, especially considering that a real-world CP dataset focused on adverse weather does not yet exist.
6 Conclusions
@@
Our experiments showed that models trained without adverse weather scenes suffer a significant drop in performance when used for inference in these conditions, whereas models trained with them still face a significant challenge. In the future, we hope to investigate the robustness of CP models to adverse weather conditions and improve upon their designs to address shortcomings arising from their use in adverse weather.
@@
As an open-source dataset, Adver-City’s open code can be seamlessly extended to include additional weather conditions such as snow or dust, new road configurations, and maps that leverage weather and density variations. Sensor configurations can also be adjusted to generate new data or enhance the realism of CARLA’s sensors. We hope that its open-source nature encourages users to share their work, extending the benefits of collaboration beyond the confines of perception algorithms.
Supplementary Material
7 Road Configurations
The placement of CAVs (Connected Autonomous Vehicles) and RSUs (Roadside Units) for each road configuration used in Adver-City is presented in Figures 6, 7, 8, 9, and 10.
The rural straight non-junction configuration (Figure 6) accounts for the longest path to be driven by the ego vehicle, creating the longest scenarios of the dataset. The ego’s path goes through a yielding intersection, then to a straight surrounded by crops on both sides, reaching an intersection with a stop sign and then proceeding to its destination. Car 1 starts closer to the ego’s destination but proceeds towards the ego’s starting position, while Car 2 starts ahead of the ego and goes to the same direction as it. RSU 1 is positioned on top of a Stop sign, while RSU 2 is positioned on a windmill tower at a higher elevation.
Figure 6: Rural straight non-junction configuration.
In the rural curved non-junction configuration (Figure 7), the ego vehicle goes through a curved road layout that first goes uphill and then downhill and is surrounded by trees on both sides. Car 1 begins the scenario ahead of the ego, going towards the same destination, while Car 2 starts closer to the ego’s destination and proceeds towards the ego’s starting position. Both RSUs are placed next to the road at ground height, simulating mobile RSUs commonly used by law enforcement.
Figure 7: Rural curved non-junction configuration.
The rural intersection configuration (Figure 8) has the shortest route to be taken by the ego vehicle, in which it approaches the intersection through a short straight with trees to its right and an open field to its left, waits for the traffic light to turn green, and then makes a left turn, briefly going through a section with buildings to its right and the open field to the left. Car 1 approaches the intersection from the West, while Car 2 approaches it from the North, both of them waiting for it to turn green until the end of the scenario. RSU 1 is positioned on top of the traffic light, while RSU 2 is positioned on top of an adjacent intersection’s traffic light, providing a unique viewpoint to the agents.
Figure 8: Rural intersection configuration.
In the urban non-junction configuration (Figure 9), the ego vehicle follows a straight road with a median strip, with tall walls to the right and houses to the left, performing a lane change shortly after the start of the scenario and then following a straight line until reaching its destination. Car 1 is positioned on the opposing traffic lane, moving towards the ego’s initial position, while Car 2 is slightly ahead of the ego, moving in the same direction as it, but starting the scenario on the lane immediately to the left of the ego. RSU 1 is positioned atop the tall wall to the right of the road, viewing the non-junction at an angle, while RSU 2 is positioned on top of a bus stop located at the side of the road.
Figure 9: Urban non-junction configuration.
The urban intersection configuration (Figure 10) is the one with the greatest number of objects, having the ego approach an intersection through a short two-way road, waiting for the traffic light to turn green, making a left turn towards a residential road and proceeding briefly through it. Car 1 is positioned on the residential road, approaching the intersection from the East, and Car 2 is positioned on the road West of the intersection, approaching it from that direction, both of them waiting by the intersection for the traffic light to turn green until the end of the scenario. Both RSUs are positioned on traffic lights at the intersection, offering a surrounding view of it.
Figure 10: Urban intersection configuration.
8 Comparison with Real-Life Datasets
Based on the t-SNE plot presented in Figure 5, we replaced each point with their corresponding sample to better visualize and interpret the data, with the new plot presented in Figure 11.
Figure 11: t-SNE plot with points replaced by their corresponding dataset samples.
A distinguishing characteristic of the Boreas dataset [4], based on the plot, is that most samples have rain splatter on the camera lens, a feature that can only be observed in a few samples of the NuScenes dataset [5].
At the top of the plot, a small cluster of Adver-City samples was separated from the rest of the dataset due to its similarity to samples from the BDD100K dataset [44], this being the presence of a white van in front of the ego vehicle.
Looking at the main Adver-City cluster, it is possible to note similar characteristics between its samples and those of the other datasets that were positioned adjacent to it. NuScenes samples positioned closer to our dataset’s main cluster have as a distinguishing characteristic a cloudy gray sky similar to those of the simulated dataset. Samples from the BDD100K dataset with high vehicle density were positioned to the left of the plot closer to samples from our dataset that also feature high vehicle density. Then, looking at the vicinity between the Adver-City and the Boreas clusters, we can observe that all the samples are characterized by having a wet pavement.